# Bivariate phase-type distributions for experience rating in disability insurance

### v2

@@ Abstract.
In this paper, we consider the problem of experience rating within the classic Markov chain life insurance framework. We begin by investigating various multivariate mixed Poisson models with mixing distributions encompassing independent Gamma, hierarchical Gamma, and multivariate phase-type. In particular, we demonstrate how maximum likelihood estimation for these proposed models can be performed using expectation-maximization algorithms, which might be of independent interest. Subsequently, we establish a link between mixed Poisson distributions and the problem of pricing group disability insurance contracts that exhibit heterogeneity. We focus on shrinkage estimation of disability and recovery rates, taking into account sampling effects such as right-censoring. Finally, we showcase the practicality of these shrinkage estimators through a numerical study based on simulated yet realistic insurance data. Our findings highlight that by allowing for dependency between latent group effects, estimates of recovery and disability rates mutually improve, leading to enhanced predictive performance.
@@ 1. Introduction
This paper builds on and extends the work of Furrer, (2019) on group shrinkage estimation of transition rates in Markov chains by allowing for dependence between latent group effects across transitions. Following closely Yslas, (2021), but focusing on mixed Poisson regressions rather than frailty models, we study multivariate phase-type distributions as a tractable and flexible class of priors for the latent group effects. The connection to Markov chain models is provided by the classic connection between Poisson regression and parametric inference in multi-state models. The phase-type approach is compared with other alternatives, including the hierarchical system of finite Gamma mixtures proposed in Norberg, (1989) as well as some naïve choices, through an in-depth simulation study related to experience rating for disability insurance. In general, the phase-type approach outperforms the alternatives.
@@
In its univariate form, a phase-type distribution is defined as the distribution of the time until absorption of a time-homogeneous, finite state-space Markov jump process with a single absorbing state, while the remaining states are transient. These distributions have found several applications in applied probability due to closed-form formulas for different functionals and, more importantly, their denseness on the set of distributions on the positive real line. In actuarial science, we find applications towards areas such as risk theory (Bladt,, 2005), loss severity modeling (Bladt and Yslas,, 2023), mortality modeling (Albrecher et al., 2022a, ), and multi-state modeling (Ahmad et al.,, 2023; Ahmad and Bladt,, 2023). Several multivariate extensions of phase-type distributions exist in the literature. Notable among them are the general MPH* class introduced in Kulkarni, (1989) and its more computationally amenable subclasses: the feed-forward type class, confer, for example, with Albrecher et al., 2022b , and the mPH class recently detailed in Bladt, (2023), which also includes applications towards severity modeling. In this work, we opt as in Yslas, (2021) for the feed-forward type class due to its mathematical convenience and its denseness on the set of the distributions on the positive orthant.
@@
Both classes of priors are compared to some naïve alternatives through an in-depth simulation study on group disability insurance. This numerical study serves a dual purpose: It demonstrates that our new specifications can be implemented smoothly for near real-word scenarios and it provides insight into the advantages and disadvantages of each specification. The phase-type specification exhibits as good or strictly better predictive performance across all relevant cases.
@@
For the application to group disability insurance, we draw on the work of Furrer, (2019), which employed an empirical Bayes approach to group experience rating in Markov chain models with the latent effects introduced on the level of transition rates. Although Furrer, (2019) provides details on the sufficiency of summary statistics such as occurrences and exposures, hereby replicating the by now well-known link between likelihood based inference for multi-state and Poisson models, the proposed shrinkage estimators are derived under a strict assumption of independence between the latent group effects. However, dependence between latent effects is to be expected, especially if only few covariates are included. Thus, allowing for dependence could enhance predictive accuracy. Concordantly, in Furrer, (2019) such an extension is highlighted as a promising avenue for future research. In light of this, the present work may be viewed as picking up exactly where Furrer, (2019) left, showcasing the potential of a more sophisticated modeling approach.
@@
The remainder of the paper is organized as follows. In Section 2, we provide an overview of mixed Poisson regression and thoughtfully analyze the model’s properties and its estimation when the mixing distribution is one of three distinct mixing distributions: independent Gamma, hierarchical Gamma, and multivariate phase-type. In Section 3, we describe the problem of experience rating for disability insurance based on Markov chain modeling and recall the connection to mixed Poisson regression. Section 4 is dedicated to a detailed numerical study based on simulated data showcasing the practical applicability of the presented models as well as the overall preferability of the phase-type specification. Finally, Section 5 concludes.
@@ 2.3. Hierarchical Gamma mixing
$\Theta_{j}\mid\Theta_{0}\sim\text{Gamma}(\Theta_{0},\delta),\quad j=1,\dots,q,\quad\delta>0.$
@@
for all $j,k=1,\dots,q$, $j\neq k$. In particular, from above, we note that the hierarchical framework allows only for positively correlated mixing variables.
@@
(2.3) $\Theta_{j}\mid\Theta_{0},\boldsymbol{Y}=\boldsymbol{y}\sim\text{Gamma}\Big{(}y_{\bullet j}+\Theta_{0},e_{\bullet j}+\frac{\eta}{\nu}\Big{)}.$
@@
The scaled mixing weights $\bar{w}_{m}$ satisfying $\sum_{m}\bar{w}_{m}=1$ can then be computed straightforwardly from the unscaled mixing weigths $w_{m}$. This interpretation in terms of conjugate priors is particularly useful for the derivation of the estimation procedure presented next. Practical experiences indicate, unfortunately, that this recurrence scheme might come with computational performance issues for large datasets.
@@ 2.3.1. Estimation
E-step. The first step is to compute the conditional expectation of the loglikelihood given the observed data. To this end, let us denote by $\boldsymbol{\beta}^{(k)}$, $\eta^{(k)}$, and $\nu^{(k)}$ the current parameters after $k$ iterations. Then, for the $(k+1)$’th iteration, the problem reduces to computing $\mathbb{E}^{(k+1)}(\Theta_{0}\mid\tilde{\boldsymbol{y}})$, $\mathbb{E}^{(k+1)}(\log(\Theta_{0})\mid\tilde{\boldsymbol{y}})$, and $\mathbb{E}^{(k+1)}(\Theta_{j}\mid\tilde{\boldsymbol{y}})$, $j=1,\dots,q$, where $\mathbb{E}^{(k+1)}(\cdot)$ denotes the expectation computed using the current parameters. We perform the calculations for one (generic) data point $\boldsymbol{y}=(\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{q})$, given that for a sample of size $N$, we simply require to sum over all data points on those expressions. First, we note that (2.3) yields
@@
$\frac{\partial}{\partial\eta}l_{c}(\boldsymbol{\beta},\nu,\eta;\tilde{\boldsymbol{y}},\tilde{\boldsymbol{x}})$ $\propto\frac{1}{q}\Big{(}\log(\nu)-\Psi(\eta)+\frac{1}{N}\sum_{n=1}^{N}\log(\theta_{0}^{(n)})\Big{)}$
@@
To compute the update parameter $\eta^{(k+1)}$, we are in principle required to find the solution to $\frac{\partial}{\partial\eta}l_{c}=0$. This equation does not possess a closed-form solution, so the use of some numerical optimizer is necessitated. However, we resort to just performing a single Netwon-Raphson step, which seems to perform rather well in practice. It is given by
@@ 2.4. Bivariate phase-type mixing
[…]
@@
with $\bm{\bm{I}}$ the identity matrix of appropriate dimension. Besides their mathematical tractability, phase-type distributions are particularly appealing as modeling tools due to their denseness within the set of distributions on the positive real line. We refer to Bladt and Nielsen, (2017) for a comprehensive account of phase-type distributions.
@@
where $\boldsymbol{\eta}$ is a vector of initial probabilities of dimension $p_{1}$, $\bm{\bm{T}}_{11}$ and $\bm{\bm{T}}_{22}$ are sub-intensity matrices of dimensions $p_{1}\times p_{1}$ and $p_{2}\times p_{2}$, respectively, and $\bm{\bm{T}}_{12}$ is a $p_{1}\times p_{2}$ matrix satisfying $\bm{\bm{T}}_{11}{\mathbf{1}}=-\bm{\bm{T}}_{12}{\mathbf{1}}$. The motivation for considering this class is that it is dense on the set of distributions on $\mathbb{R}_{+}^{2}$ (see Albrecher et al., 2022b, ), thus allowing for the modeling of any dependence structure. Moreover, this class can be estimated via an EM algorithm (see Ahlström et al.,, 1999). The more recently introduced class of multivariate phase-type distributions in Bladt, (2023), named the mPH class, possesses similar properties, and it is, in fact, a subfamily of the distributions with joint density (2.6). However, in the present work we focus only on the class described by (2.6) to keep the presentation concise.
@@ 2.4.1. Estimation
E-step. We denote by $\boldsymbol{\beta}^{(k)}$, $\boldsymbol{\eta}^{(k)}$, $\bm{\bm{T}}_{11}^{(k)}$, $\bm{\bm{T}}_{12}^{(k)}$, and $\bm{\bm{T}}_{22}^{(k)}$ the current parameters after $k$ iterations. Then, for the $(k+1)$’th iteration, we are required to compute $\mathbb{E}^{(k+1)}(\Theta_{1}\mid\tilde{\boldsymbol{y}})$, $\mathbb{E}^{(k+1)}(\Theta_{2}\mid\tilde{\boldsymbol{y}})$, and $\mathbb{E}^{(k+1)}(\log(f_{\bm{\bm{\Theta}}}(\bm{\bm{\Theta}}))\mid\tilde{\boldsymbol{y}})$. As before, it is enough to perform the calculations for one (generic) data point $\boldsymbol{y}=(\boldsymbol{y}_{1},\boldsymbol{y}_{2})$. Note that (2.7) yields
@@
A general closed-form solution to the integral above is not available. However, we only require maximization of this expression with respect to the bivariate phase-type parameters, which can be carried out numerically, as we discuss next.
@@
$\mathbb{E}^{(k+1)}(\log(f_{\bm{\bm{\Theta}}}(\bm{\bm{\Theta}};\boldsymbol{\eta},\bm{\bm{T}}_{11},\bm{\bm{T}}_{12},\bm{\bm{T}}_{22}))\mid\tilde{\boldsymbol{y}})$
@@
$\frac{1}{N}\sum_{n=1}^{N}\int_{\mathbb{R}_{+}^{2}}\log(f_{\bm{\bm{\Theta}}}(\boldsymbol{\theta};\boldsymbol{\eta},\bm{\bm{T}}_{11},\bm{\bm{T}}_{12},\bm{\bm{T}}_{22}))f_{\bm{\bm{\Theta}}|\boldsymbol{Y}}(\boldsymbol{\theta}|\boldsymbol{y}^{(n)})\,\mathrm{d}\boldsymbol{\theta}$
@@
is a joint density function. Thus, the problem translates into fitting a bivariate phase-type distribution to (2.9). In Albrecher et al., 2022b , it is shown how the EM algorithm in Ahlström et al., (1999) can be modified for such a purpose. We also refer to Yslas, (2021) for further details.

### v3

@@ Abstract.
In this paper, we consider the problem of experience rating within the classic Markov chain life insurance framework. We begin by establishing a link between mixed Poisson distributions and the problem of pricing group disability insurance contracts that exhibit heterogeneity. We focus on shrinkage estimation of disability and recovery rates, taking into account sampling effects such as right-censoring. We then investigate some specific multivariate mixed Poisson models with mixing distributions encompassing independent Gamma, hierarchical Gamma, and multivariate phase-type. In particular, we demonstrate how maximum likelihood estimation for these models can be performed using expectation-maximization algorithms, which might be of independent interest. Finally, we showcase the practicality of the proposed shrinkage estimators through a numerical study based on simulated yet realistic insurance data. Our findings highlight that by allowing for dependency between latent group effects, estimates of recovery and disability rates mutually improve, leading to enhanced predictive performance.
@@ 1. Introduction
This paper builds on and extends the work of Furrer, (2019) on group shrinkage estimation of transition rates in Markov chains by allowing for dependence between latent group effects across transitions. Following closely Yslas, (2025), but focusing on mixed Poisson regressions rather than frailty models, we study multivariate phase-type distributions as a tractable and flexible class of priors for the latent group effects. The connection to Markov chain models is provided by the classic connection between Poisson regression and parametric inference in multi-state models. The phase-type approach is compared with other alternatives, including the hierarchical system of finite Gamma mixtures proposed in Norberg, (1989) as well as some naïve choices, through an in-depth simulation study related to experience rating for disability insurance. Overall, the phase-type approach outperforms the alternatives.
@@
In its univariate form, a phase-type distribution is defined as the distribution of the time until absorption of a time-homogeneous, finite state-space Markov jump process with a single absorbing state, while the remaining states are transient. These distributions have found several applications in applied probability due to closed-form formulas for different functionals and, more importantly, their denseness on the set of distributions on the positive real line. In actuarial science, we find applications towards areas such as risk theory (Bladt,, 2005), loss severity modeling (Bladt and Yslas,, 2023), mortality modeling (Albrecher et al., 2022a, ), and multi-state modeling (Ahmad et al.,, 2023; Ahmad and Bladt,, 2024). Several multivariate extensions of phase-type distributions exist in the literature. Notable among them are the general MPH* class introduced in Kulkarni, (1989) and its more computationally amenable subclasses: the feed-forward type class, confer, for example, with Albrecher et al., 2022b , and the mPH class recently detailed in Bladt, (2023), which also includes applications towards severity modeling. In this work, we opt as in Yslas, (2025) for the feed-forward type class due to its mathematical convenience and its denseness on the set of the distributions on the positive orthant.
@@
Both classes of priors are compared to some naïve alternatives through an in-depth simulation study on group disability insurance. For disability insurance, there is often a natural group structure, especially if the insurance is provided through a corporate (pension) scheme, in which case each company constitutes a more or less closed group. Due to differences in characteristics across companies, including differences in internal practices, there is often substantial heterogeneity between groups – also compared to the heterogeneity within groups – which in turn gives rise to non-negligible group effects. The numerical study serves a dual purpose: It demonstrates that our new specifications can be implemented smoothly for near real-word scenarios and it provides insight into the advantages and disadvantages of each specification. The phase-type specification essentially exhibits as good or strictly better predictive performance across all relevant cases.
@@
For the application to group disability insurance, we draw on the work of Furrer, (2019), which employed an empirical Bayes approach to group experience rating in Markov chain models with the latent effects introduced on the level of transition rates. Linear Bayes methodology, rather than empirical Bayes, constitutes an alternative which in actuarial science is often referred to as credibility theory; for an overview, see Bühlmann and Gisler, (2005). Although Furrer, (2019) provides details on the sufficiency of summary statistics such as occurrences and exposures, hereby replicating the by now well-known link between likelihood based inference for multi-state and Poisson models, the proposed shrinkage estimators are derived under a strict assumption of independence between the latent group effects. However, dependence between latent effects is to be expected, especially if only few covariates are included. Thus, allowing for dependence could enhance predictive accuracy. Concordantly, in Furrer, (2019) such an extension is highlighted as a promising avenue for future research. In light of this, the present work may be viewed as picking up exactly where Furrer, (2019) left, showcasing the potential of a more sophisticated modeling approach.
@@
The remainder of the paper is organized as follows. In Section 2, we describe the problem of experience rating for disability insurance based on Markov chain modeling and recall the connection to mixed Poisson regression. In Section 3, we provide an overview of mixed Poisson regression and thoughtfully analyze the model’s properties and its estimation when the mixing distribution is one of three distinct mixing distributions: independent Gamma, hierarchical Gamma, and multivariate phase-type. Section 4 is dedicated to a detailed numerical study based on simulated data showcasing the practical applicability of the presented models as well as the overall preferability of the phase-type specification. Finally, Section 5 concludes.
2. Experience rating for disability insurance
The area of application of this paper is experience rating in disability insurance; that is, we seek to price group disability insurance contracts subject to heterogeneity. Subsection 2.1 contains a description of the insurance coverage of interest. In Subsection 2.2, we consider shrinkage estimation of disability and recovery rates subject to, among other things, right-censoring and following closely Furrer, (2019). In particular, we provide a link to mixed Poisson models, which are the focal point of Section 3.
2.1. Disability coverage
$U(t)=t-\sup\{s\in[0,t]:Z_{s}\neq Z_{t}\}.$
Finally, $\varepsilon\in[0,\infty)$ is a waiting period, $\tau\in(0,\infty]$ is the coverage period, $t\mapsto b_{ai}(t)$ are deterministic benefits upon disability, and $t\mapsto b_{i}(t)$ is a deterministic benefit rate during disability. The interpretation of (2.1) is as follows. If the insured becomes disabled in the coverage period, say at time $t$ with $t\leq\tau$, then they receive at exactly time $t$ the payment $b_{ai}(t)$. Additionally, they receive a payment rate of $s\mapsto b_{i}(s)$ from time $t+\varepsilon$ until recovery or death, whichever occurs first; should it happen that they recover or die before time $t+\varepsilon$, they receive however only the payment $b_{ai}(t)$.
Example 2.1.
In the following, we suppose that the payment functions $b_{ai}$ and $b_{i}$ are suitably regular, for instance piecewise continuous, and that there exists a maximal contract time $T\in(0,\infty)$ such that $b_{ai}(t)=b_{i}(t)=0$ for $t>T$. This is always the case in practice (with $T$ related to, for instance, the legal retirement age). Furthermore, we suppose that the initial state of $Z$ is $a$, and we naturally take $d$ to be absorbing.
We further assume that $Z$ is Markovian and admits transition rates $t\mapsto\mu_{jk}(t)$, $\in\mathcal{S}, jkj\neq k$, such that
$\mu_{jk}(t)=\lim_{h\downarrow 0}\frac{\mathbb{P}(Z_{t+h}=k\,|\,Z_{t}=j)}{h}$
In other words, we have
The model is illustrated in Figure 2.1.
[…]
where $t\mapsto r(t)$ is a suitably regular, say piecewise continuous, interest rate of choice. We are, of course, implicitly requiring that $s\leq t$. The following result is a consequence of, for instance, Corollary 7.2 in Adékambi and Christiansen, (2017).
Proposition 2.2.
It holds almost everywhere that
Section 8 in Adékambi and Christiansen, (2017) provides a numerical scheme for the prospective reserves based on (2.2) in conjunction with the obvious boundary conditions $V_{a}(T)=0$ and $V_{i}(T,s)=0$, the system given by (2.2) being a special case of Thiele’s differential (or integral) equations for semi-Markov models. The interpretation of (2.2) is similar to the classic case with deterministic sojourn payment rates and transition payments, confer with Hoem, (1969).
Remark 2.1.
This leads to a particularly straightforward numerical scheme.
2.2. Shrinkage estimation
Experience rating for group disability insurance is closely related to shrinkage estimation of group effects. Experience rating concerns the improvement of a portfolio level assessment through the use of information about a specific (sub-)group’s performance. Shrinkage estimators improve, either explicitly or implicitly, a naïve estimate by bringing it closer to a value supplied by additional information. If we think of the raw group effect as the naïve estimate and the portfolio level as the value supplied by additional information, we see that shrinkage estimation and experience rating seek the same goal, be it from different directions.
In this context, Furrer, (2019) suggests estimation of the transition rates using empirical Bayes methods. Disregarding mortality, the ‘simple shrinkage’ estimation of Furrer, (2019) is based on the following structural assumptions:
$O_{ai}^{k}\,|\,(\Theta_{ai},\Theta_{ia}){=(\theta_{ai},\theta_{ia})}$ $\sim\text{Poisson}\big({\theta_{ai}}E_{a}^{k}\mu_{ai}(t_{k})\big),$
$O_{ia}^{k}\,|\,(\Theta_{ai},\Theta_{ia}){=(\theta_{ai},\theta_{ia})}$ $\sim\text{Poisson}\big({\theta_{ia}}E_{i}^{k}\mu_{ia}(t_{k})\big).$
This provides a direct link to Section 3.
The independence assumption is disadvantageous, as in practice you typically observe substantial dependence between the group effects, especially if the group effects indirectly capture differences in the types of disabilities experienced by individuals across groups. Stress related claims are, for example typically more frequent but also shorter in duration than claims stemming from diseases of or injuries to the musculoskeletal system. This entails positive correlation between the disability and recovery rate. In light hereof, the next section contains a theoretical investigation of mixed Poisson models with other types of mixing than that of (2.5), namely hierarchical and phase-type mixings. This theoretical investigation is finally supplemented by a numerical study devoted to experience rating for disability insurance. By allowing for dependence between latent group effects, estimation of the recovery rate may borrow strength from the estimation of the disability rate – and vice versa. Everything else being equal, this should lead to improvements in predictive performance.
@@ 3.3. Hierarchical Gamma mixing
$\Theta_{j}\mid\Theta_{0}{=\theta_{0}}\sim\text{Gamma}({\theta_{0}},\delta),\quad j=1,\dots,q,\quad\delta>0.$
@@
for all $1,\dots,q, jkj\neq k$. In particular, from above, we note that the hierarchical framework allows only for positively correlated mixing variables. Furthermore, the variances of the marginal distributions cannot differ. Therefore, even in the limiting case $\eta\to\infty,\nu\to\infty,\eta/\nu\to\delta$, we do not recover the previous independent Gamma mixing.
@@
(3.3) $\Theta_{j}\mid\Theta_{0}{=\theta_{0}},\boldsymbol{Y}=\boldsymbol{y}\sim\text{Gamma}\Big(y_{\bullet j}+{\theta_{0}},e_{\bullet j}+\frac{\eta}{\nu}\Big).$
@@
The scaled mixing weights $\bar{w}_{m}$ satisfying $\sum_{m}\bar{w}_{m}=1$ can then be computed straightforwardly from the unscaled mixing weigths $w_{m}$. This interpretation in terms of conjugate priors is particularly useful for the derivation of the estimation procedure presented next. Practical experiences indicate, unfortunately, that this recurrence scheme might lead to arithmetic overflow for large datasets.
@@ 3.3.1. Estimation
[…]
@@
[…]
@@
To compute the update parameter $\eta^{(k+1)}$, we are in principle required to find the solution to $\frac{\partial}{\partial\eta}l_{c}=0$. This equation does not possess a closed-form solution, so the use of some numerical optimizer is necessitated. We propose performing a single Newton-Raphson step, which is both computationally efficient and sufficient to ensure convergence of the ECM algorithm (confer with Meng and Rubin,, 1993). Thus, the update is given by
@@ 3.4. Bivariate phase-type mixing
where $\bm{\bm{T}}$ is a $p\times p$ sub-intensity matrix and $\boldsymbol{t}$ is a $p$-dimensional column vector satisfying $\boldsymbol{t}=-\bm{\bm{T}}\,{\mathbf{1}}$ with $\mathbf{1}$} the $p$-dimensional column vector of ones. Assume that the process starts in a transit state $k$ with probability $\pi_{k}={\mathbb{P}}(J_{0}=k)$, $k=1,\dots,p$, and that $\sum_{k=1}^{p}\pi_{k}=1$, that is, the process cannot start in the absorbing state. Finally, let $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{p})$ be the vector of initial probabilities. Then, we say that the time until absorption $\Theta=\inf\{t\geq 0:J_{t}=p+1\}$ has a phase-type distribution with representation $(\boldsymbol{\pi},\bm{\bm{T}})$, and we write $\Theta\sim\mbox{PH}(\boldsymbol{\pi},\bm{\bm{T}})$.
@@
with $\bm{\bm{I}}$ the identity matrix of appropriate dimension. Besides their mathematical tractability, phase-type distributions are particularly appealing as modeling tools due to their denseness within the set of distributions on the positive real line. Notably, they also include several important distributional families as special cases, such as Erlang distributions (that is Gamma distributions with integer shape parameters), general Coxian laws, and mixtures thereof. We refer to Bladt and Nielsen, (2017) for a comprehensive account of phase-type distributions.
@@
Remark 3.2.
We note that one can alternatively employ other phase-type extensions as mixing distributions. In particular, certain inhomogeneous phase-type specifications introduced in Albrecher and Bladt, (2019) admit closed-form Laplace transforms and can be extended to the multivariate setting via the transformation approach in Albrecher et al., 2022b . While such models are natural candidates, we focus here on the conventional phase-type case, as it provides the most direct extension of the independent Gamma mixing considered in Furrer, (2019), and already offers sufficient flexibility for our application.
@@
where $\boldsymbol{\eta}$ is a vector of initial probabilities of dimension $p_{1}$, $\bm{\bm{T}}_{11}$ and $\bm{\bm{T}}_{22}$ are sub-intensity matrices of dimensions $p_{1}\times p_{1}$ and $p_{2}\times p_{2}$, respectively, and $\bm{\bm{T}}_{12}$ is a $p_{1}\times p_{2}$ matrix satisfying $\bm{\bm{T}}_{11}{\mathbf{1}}=-\bm{\bm{T}}_{12}{\mathbf{1}}$. The motivation for considering this class is that it is dense on the set of distributions on $\mathbb{R}_{+}^{2}$ (see Albrecher et al., 2022b, ), thus allowing for the modeling of any dependence structure. Some instances of bivariate specifications that fall under this framework include the case of independent phase-type components (in particular, independent Erlangs) and multivariate Erlang distributions constructed via mixtures of independent Erlangs. Moreover, this class can be estimated via an EM algorithm (see Ahlström et al.,, 1999). The more recently introduced class of multivariate phase-type distributions in Bladt, (2023), named the mPH class, possesses similar properties, and it is, in fact, a subfamily of the distributions with joint density (3.6). Extensions beyond the bivariate case are possible, either through the construction outlined in Albrecher et al., (2021) or via the mPH class, although such generalizations come with additional computational challenges (confer with Yslas,, 2025, for a detailed discussion). In the present work, however, we focus only on the bivariate class described by (3.6) to keep the presentation concise and in accordance with the present application.
@@ 3.4.1. Estimation
[…]
@@
A general closed-form solution to the integral above is not available. However, an explicit formula is not required, as the subsequent maximization of this expression with respect to the phase-type parameters can be translated into fitting a bivariate phase-type distribution to a known target distribution, as earlier noted in Yslas, (2025). This approach eliminates the need for a closed-form solution, allowing the maximization to be carried out numerically, as discussed next.
@@
[…]
@@
[…]
$=\int_{\mathbb{R}_{+}^{2}}\log(f_{\bm{\bm{\Theta}}}(\boldsymbol{\theta};\boldsymbol{\eta},\bm{\bm{T}}_{11},\bm{\bm{T}}_{12},\bm{\bm{T}}_{22}))\left[\frac{1}{N}\sum_{n=1}^{N}f_{\bm{\bm{\Theta}}|\boldsymbol{Y}}(\boldsymbol{\theta}|\boldsymbol{y}^{(n)})\right]\,\mathrm{d}\boldsymbol{\theta}$
@@
is a joint density function. Finally, note that maximizing the above expression with respect to the phase-type parameters is equivalent to minimizing the Kullback–Leibler divergence between the bivariate phase-type distribution and (3.9). The latter, in turn, can be carried out by fitting a bivariate phase-type model to the target distribution specified by (3.9). In Albrecher et al., 2022b , it is shown how the EM algorithm in Ahlström et al., (1999) can be modified for such a purpose. We also refer to Yslas, (2025) for further details.